# On the Design and Analysis of LLM-Based Algorithms

### v1

@@ 3.6 Example: retrieval-augmented generation
In our last example, we consider a multiple-needle generalization of the previous retrieval task, which can be regarded as a synthetic and simplified version of retrieval-augmented generation (RAG) [LPP+20, GXG+24]. In particular, we consider fine-grained sentence-level retrieval by LLMs, rather than document-level or chunk-level retrieval by certain similarity measure of dense embedding vectors; in other words, we skip the embedding steps in practical RAG.
More concretely, we let the input text be composed of sentences of the form “The {i}-th digit of the passcode to the {colored object}” is {digit}”, where $i\in[6]$. The algorithm is asked to answer the question “What is the 6-digit passcode to the {targeted object}?”. Compared with the previous single-needle retrieval task, here the algorithm need to retrieve multiple needles, each for one digit of the targeted object, in order to answer the question correctly; moreover, the final aggregation step requires certain capability of reasoning or summarization over the retrieved needles.

### v2

@@ 3.6 Example: retrieval-augmented generation
Our next example is a multiple-needle generalization of the previous retrieval task, which can be regarded as a synthetic and simplified version of retrieval-augmented generation (RAG) [LPP+20, GXG+24]. In particular, we consider fine-grained sentence-level retrieval by LLMs, rather than document-level or chunk-level retrieval by certain similarity measure of dense embedding vectors.
More concretely, suppose that the input text is composed of sentences of the form “The {i}-th digit of the passcode to the {colored object}” is {digit}”, where $i\in[6]$. The algorithm is asked to answer the question “What is the 6-digit passcode to the {targeted object}?”. Compared with the previous single-needle retrieval task, here the algorithm need to retrieve multiple needles, each for one digit of the targeted object, in order to answer the question correctly; moreover, the final aggregation step requires certain capability of reasoning or summarization over the retrieved needles.